OpenStudy (anonymous):

You attend a baseball game with friends and visit the snack stand between innings. Soda costs $2 a can, and a box of popcorn costs $1.50. You have $15 with you, and you wish to buy more cans of soda than boxes of popcorn. x represents the number of cans of soda and y the number of boxes of popcorn, how will this be shown on a graph? Please help

OpenStudy (anonymous):

u want to buy 8 dollares 50cents of soda and the rest on popcorn

OpenStudy (anonymous):

\[2x+1.5y \le15\] \[x>y\]

OpenStudy (anonymous):

plot the equality \[2x+1.5y=15\] Should be a straight line, then find the feasible region for the inequality \[2x+1.5y<=15\] by plugging in values of x and y such that the point (x,y) does not lie on the line 2x+1.5y=15 If the inequality is true, the feasible region is towards the point (x,y) , otherwise opposite side